OpenStudy (anonymous):

If sin a = .25 and cos b = .25, which of the following is sin b/2 + cos a/2 A. 1.60 B. 1.32 C. .26 D. 1.06

OpenStudy (whpalmer4):

Your problem doesn't make sense as stated...

OpenStudy (anonymous):

@whpalmer4 If sin a = .25 and cos b = .25, which of the following is sin b/2 + cos a/2

OpenStudy (whpalmer4):

I don't see anything following...

OpenStudy (anonymous):

I just fixed the problem in the box. @whpalmer4

OpenStudy (whpalmer4):

"which of the following is sin b/2 + cos a/2" — there's nothing following!

OpenStudy (anonymous):

@whpalmer4 A. 1.60 B. 1.32 C. .26 D. 1.06

OpenStudy (whpalmer4):

Do you know the half-angle identities for sin and cos?

OpenStudy (whpalmer4):

\[\sin^2 u = \frac{1-\cos 2u}{2}\]\[\sin u = \sqrt{\frac{1-\cos 2u}{2}}\] \[\cos^2 u = \frac{1+\cos 2u}{2}\]\[\cos u = \sqrt{\frac{1+\cos 2u}{2}}\]

OpenStudy (whpalmer4):

Next, we draw the two triangles representing \(\sin a = 0.25\) and \(\cos b = 0.25\)  We'll use the triangles to find the value of \(\cos a\) and \(\sin b\) if needed.

OpenStudy (whpalmer4):

We want \(\sin(b/2) + \cos(a/2)\) which we can construct with the identities given above if we let \(u = 0.25/2\). Then we'll have \[\sin(b/2) + \cos(a/2) = \sqrt{\frac{1-\frac{1}{4}}{2}} + \sqrt{\frac{1+\cos(a)}{2}} = \sqrt{\frac{3}{8}} + \sqrt{\frac{1+\cos(a)}{2}}\] Looking at that triangle we drew, we see that \(\cos(a) = \dfrac{\sqrt{15}}{4}\) so our equation becomes \[\sin(b/2) + \cos(a/2) =\sqrt{\frac{3}{8}} + \sqrt{\frac{1+\frac{\sqrt{15}}{4}}{2}} = \sqrt{\frac{3}{8}}+ \sqrt{\frac{4+\sqrt{15}}{8}}\] Punch that into your calculator and compare with your answer choices...